[Hallc_running] RC meeting today at 1615. Meeting ID : 861439920

Jay Benesch benesch at jlab.org
Sun Mar 22 12:41:08 EDT 2020

There will be an RC meeting at 1615 today via BlueJeans.  861439920

Work is underway in the hall to button everything up.  Dave Gaskell 
estimated 2-3 hours at noon, call it 1500.  Ops will have to sweep and 
bring MBSY3C up slowly.  They will need to check vertical RF separation. 
  The amps have been on all week so the only adjustment needed might be 
a small phase change.  Say 1800 before Ops is ready to pull the dumplette.

Given https://logbooks.jlab.org/entry/3807201 the Shift leader will have 
to work closely with Ops to define different lock positions for each of 
the targets.

Brad has promised me a detailed run plan after the hall is locked up.

There was little of note at the 9am PD meeting. IN response to questions 
from Stuart, Dave Hamlette estimated one to two days to secure the Hall 
A calcium targets, mostly solo work by Dave Meekins.  Jonathan Creel 
estimated 3-4 days to bring the cryo system to stability at 4K with 
minimum helium loss. Most of this can be done remotely.  Jonathan 
emphasized that cryo personnel will have to be on site daily to meet LN2 
trucks and walk all the systems to look for problems with the rotating 
machinery that aren't visible in EPICS.  The PD meeting time will return 
to 0800 Monday.

More information about the Hallc_running mailing list